# EarthquakeNPP: A Benchmark for Earthquake Forecasting with Neural Point Processes

### v2

@@ Abstract
For decades, classical point process models, such as the epidemic-type aftershock sequence (ETAS) model, have been widely used for forecasting the event times and locations of earthquakes. Recent advances have led to Neural Point Processes (NPPs), which promise greater flexibility and improvements over such classical models. However, the currently-used benchmark for NPPs does not represent an up-to-date challenge in the seismological community, since it contains data leakage and omits the largest earthquake sequence from the region. Additionally, initial earthquake forecasting benchmarks fail to compare NPPs with state-of-the-art forecasting models commonly used in seismology. To address these gaps, we introduce EarthquakeNPP: a collection of benchmark datasets to facilitate testing of NPPs on earthquake data, accompanied by an implementation of the state-of-the-art forecasting model: ETAS. The datasets cover a range of small to large target regions within California, dating from 1971 to 2021, and include different methodologies for dataset generation. Benchmarking experiments, using both log-likelihood and generative evaluation metrics widely recognised in seismology, show that none of the five NPPs tested outperform ETAS. These findings suggest that current NPP implementations are not yet suitable for practical earthquake forecasting. Nonetheless, EarthquakeNPP provides a platform to foster future collaboration between the seismology and machine learning communities.
@@ 1 Introduction
Operational earthquake forecasting by global governmental organisations such as the US Geological Survey (USGS) necessitates the development of models which can forecast the times and locations of damaging earthquakes. While model development is ongoing in the seismology community, recent improvements have relied upon refinement of a spatio-temporal point process model known as the Epidemic-Type Aftershock Sequence (ETAS) model [53, 54], despite significant growth in available data [79, 72, 62, 90, 51, 80, 50].
In contrast, the machine learning community has offered promising advancements over classical point process models like ETAS with Neural Point Process (NPP) models, showcasing greater flexibility [9, 56, 69, 31, 5, 97, 96]. While some initial benchmarking of these models has been conducted on an earthquake dataset in Japan, these experiments lack relevance for stakeholders in the seismology community. The benchmark omits the largest earthquake sequence from the region, introduces data leakage with non-sequential train-test splits, and doesn’t compare against state-of-the-art models like ETAS.
@@
Although initial benchmarking shows that none of the five tested NPP implementations outperform ETAS, EarthquakeNPP is designed to support ongoing model development and evaluation. In addition to the standard log-likelihood metric common in the NPP literature, the platform incorporates the generative evaluation procedures used in seismology for more rigorous benchmarking. This ensures that future NPPs (and other models such as time series approaches [88] and Bayesian point processes [68]) can have direct relevance to seismological stakeholders. All datasets, experiments, and documentation are available at https://github.com/ss15859/EarthquakeNPP/tree/main.
@@ 1.1 Related Work
Benchmarking by the NPP Community. Chen et al. [5] introduced an earthquake dataset for benchmarking the Neural Spatio-temporal Point Process (NSTPP) model using a global dataset from the U.S. Geological Survey, focusing on Japan from 1990 to 2020. They considered earthquakes with magnitudes above 2.5, splitting the data into month-long segments with a 7-day offset. They exclude earthquakes from November 2010 to December 2011, deeming these sequences "too long" and "outliers." However, this period includes the 2011 Tohoku earthquake [49], the largest earthquake recorded in Japan and the fourth largest in the world, at magnitude 9.0. This exclusion renders the benchmarking experiment irrelevant for seismologists, as it is precisely these large earthquakes and their aftershocks that are crucial to forecast due to their damaging impact.
@@
The dataset segments are divided for training, testing, and validation. Instead of a chronological partitioning that mirrors operational forecasting, the segments are assigned in an alternating pattern. This approach misrepresents a realistic forecasting scenario and inflates performance measures due to earthquake triggering [13] - a form of data leakage. Since the model is tested on windows immediately preceding training windows, it exploits causal dependencies backwards in time.
@@
Although earthquakes with magnitudes above 2.5 are considered by Chen et al. [5], following a change in USGS policy on global data collection, from 2009 onwards, only events above magnitude 4.0 are recorded in the dataset. For earthquake forecasting in Japan, seismologists use datasets from Japanese data centers since they are more comprehensive and complete than global datasets. Section A.2 describes the biases incurred from such data missingness.
@@
Chen et al. [5] benchmark their model against another spatio-temporal model, Neural Jump SDEs [31], and a temporal-only Hawkes process, even though a spatio-temporal Hawkes process would provide a more rigorous benchmark. Subsequent papers adopting this benchmark [97, 94, 96] similarly lack comparisons to a spatio-temporal Hawkes process, benchmarking instead against temporal-only or spatial-only baselines or other spatio-temporal NPPs.
@@
Benchmarking by the Seismology Community. Model comparison has been crucial in the development of earthquake forecasting models since their inception [34, 53]. The Collaboratory for the Study of Earthquake Predictability (CSEP) [41, 66, 64, 30] [https://cseptesting.org/ ] aims to unify the framework for earthquake model testing and evaluation, hosting retrospective and fully prospective forecasting experiments globally. CSEP benchmarks short-term models using performance metrics that require forecasts to be generated by simulating many repeat sequences over a specified time horizon (typically one day). These simulated forecasts are compared by discretizing time and space intervals, with test statistics calculated for event counts, magnitudes, locations, and times. The simulation-based approach allows the inclusion of generative models that don’t output explicit earthquake probabilities (i.e., a likelihood), and enables evaluation of the full distribution of entire sampled sequences.
@@
Two existing works benchmark NPPs for earthquake forecasting within the seismology community. The first by Dascher-Cousineau et al. [8] extends a temporal-only NPP from Shchur et al. [69] to include earthquake magnitudes. The second by Stockman et al. [77] extends another temporal-only model by Omi et al. [56] to target larger magnitude events. Both models are benchmarked against a temporal ETAS model, showing moderate improvements over the baseline. Extending these models to include spatial data is necessary for further testing and potential operational use in the seismological community.
@@ 2.1 Spatio-Temporal Point Processes
A spatio-temporal point process is a continuous-time stochastic process that models the random number of events $N(S\times(t_{a},t_{b}])$ which occur in a space-time interval $\mathcal{S}\times(t_{a},t_{b}],\ \mathcal{S}\subseteq\mathbb{R}^{2},\ (t_{a},t_{b}]\in\mathbb{R}^{+}$. This process is typically defined by a non-negative conditional intensity function
@@
Most models specify the conditional intensity function, though some [e.g. 69, 5, 94] directly model this pdf. Model parameters are typically estimated by maximizing the log-likelihood of observed events within a training time interval $T_{0},T_{1}]$ and spatial region $\mathcal{S}$,
@@
For model evaluation and comparison, the log-likelihood of observing events in the test set can be used as a performance metric. This is consistent with a wealth of literature in the seismology community [see 95, and references therein] as well as the wider general point process literature [7], which now includes neural point processes [70]. The metric evaluates models that output probability distributions over their predictions and consequently penalises models that are overconfident. Although evaluating on events in the test set, the test log-likelihood, $\log p\left((t_{i},\mathbf{x}_{i})|t_{i}\in[T_{2},T_{3}],\mathcal{H}_{T_{2}}\right)$, may still contain dependence upon events prior to the test window $T_{2},T_{3}]$, typically contained in the history $\mathcal{H}_{T_{2}}$ of the intensity function. Comparing the difference in mean log-likelihood per event provides the information gain from one model to another [7].
Point processes are the dominant modeling approach in the seismology community, used extensively in both real-time operational earthquake forecasting [47] and established benchmarking experiments (CSEP) [81, 60]. The point process representation of earthquake data aligns naturally with their occurrence as discrete events in time [33]. Furthermore, this modeling approach is favored over discretized forecasting models (e.g., time series) because it eliminates the need for optimizing binning strategies and allows for immediate updates, rather than waiting until the end of a time bin — a delay that could miss critical, potentially damaging events.
@@ 2.2 ETAS
The Epidemic Type Aftershock Sequence (ETAS) model [54] is a spatio-temporal Hawkes process which models how earthquakes cluster in time and space. It has been adopted for operational earthquake forecasting by government agencies in California [44], New-Zealand [6], Italy [74], Japan [57] and Switzerland [48], and performs consistently well in CSEP’s retrospective and fully prospective forecasting experiments [e.g. 93, 60, 81, 4, 38, 39, 40]. The general formulation of the model is
@@
where $\mu$ is a constant background rate of events, $g(\cdot,\cdot,\cdot)$ is a non-negative excitation kernel which describes how past events contribute to the likelihood of future events and $m_{i}$ are the associated magnitudes of each event. The equivalent formulation as a Hawkes branching process accompanies a causal branching structure $\mathbf{B}$. This concept broadly aligns with the understanding of the physics of earthquake triggering and interaction, e.g. via dynamic wave triggering [3] and static stress triggering [16, 39].
Although ETAS can be fit by maximizing the log-likelihood function directly, parameter estimation is typically performed by simultaneously estimating the branching structure $\mathbf{B}$. Veen and Schoenberg [87] developed an Expectation Maximisation (EM) procedure, which maximises the marginal likelihood over the unobserved branching structure, $\log\int p(\mathcal{H}_{T_{1}}|\mathbf{B},\mathbf{\theta})p(\mathbf{B}|\mathbf{\theta})d\mathbf{B}$ through the iteration
@@
Figure 1: Earthquakes contained in the observational datasets found in EarthquakeNPP. Colours indicate the respective datasets, including the target region, magnitude of completeness $M_{c}$, number of events and the time period that the dataset spans. In red is a fault map from the GEM Global Active Faults Database [78].
The formulation of the ETAS model we present in the EarthquakeNPP benchmark is implemented in the etas python package by Mizrahi et al. [46]. It defines the triggering kernel as
@@
where $r^{2}$ is the squared distance between events and $k,a,c,\omega,\tau,d,\gamma,\rho$ are the learnable parameters along with the constant background rate $\mu$. This triggering kernel is derived from statistical distributions found through decades of observational studies [84, 83, 85] and several of the learnable parameters have been linked to physical properties of the earthquake rupture process [85, 29].
@@ 3 EarthquakeNPP Datasets
The EarthquakeNPP datasets (Figure 1) encompass earthquake records, including timestamps, geographical coordinates, and magnitudes, documented within California from 1971 to 2021. California, with its dense network and high seismic hazard, has been extensively studied, demonstrating the utility of forecasting algorithms [14, 11, 12]. It encompasses the San Andreas fault plate boundary system [99] and includes modern high-resolution catalogs with numerous small magnitude earthquakes, offering potential for new, more expressive models.
Notebooks to access and preprocess these public datasets along with the associated benchmarking experiment are publicly accessible at https://github.com/ss15859/EarthquakeNPP/tree/main, accompanied by more detailed documentation for each dataset. A summary of how earthquake datasets are generated, along with the associated challenges of using earthquake catalog data can be found in Appendix A. Table 1 provides a short summary of each EarthquakeNPP dataset.
@@
Development The U.S. Geological Survey (USGS) National Earthquake Information Center (NEIC) monitors global earthquakes (Mw 4.5 or larger) and provides complete seismic monitoring of the United States for all significant earthquakes (> Mw 3.0 or felt). Its contributing seismic networks have produced the Advanced National Seismic System (ANSS) Comprehensive Catalog of Earthquake Events and Products. The Southern California Seismic Network (SCSN) has developed and maintained the standard earthquake catalog for Southern California [28] since the Caltech Seismological Laboratory began routine operations in 1932. Significant network improvements since the 1970s and 1980s reduced the catalog completeness from Mw 3.25 to Mw 1.8. White et al. [90] created an enhanced catalog focusing on the San Jacinto fault region, using a dense seismic network in Southern California. This denser network, combined with automated phase picking (STA/LTA), ensures a 99% detection rate for earthquakes greater than Mw 0.6 in a specific subregion [90]. Using data collected by the SCSN, Ross et al. [62] generated a denser catalog by reanalyzing the same waveform data with a template matching procedure that looks for cross-correlations with the wavetrains of previously detected events.
@@
forecasting (USGS utilizes ComCat in aftershock forecasts they release to the public.)
Three magnitude thresholds (Mw 2.0, 2.5, 3.0) explore the effect of truncation on forecasting model performance. To explore if newly detected low magnitude earthquakes contain additional predictive information. To explore if newly detected low magnitude earthquakes contain additional predictive information (with different detection methodology).

### v3

@@ Abstract
For decades, classical point process models, such as the epidemic-type aftershock sequence (ETAS) model, have been widely used for forecasting the event times and locations of earthquakes. Recent advances have led to Neural Point Processes (NPPs), which promise greater flexibility and improvements over such classical models. However, the currently-used benchmark for NPPs does not represent an up-to-date challenge in the seismological community, since it contains data leakage and omits the largest earthquake sequence from the region. Additionally, initial earthquake forecasting benchmarks fail to compare NPPs with state-of-the-art forecasting models commonly used in seismology. To address these gaps, we introduce EarthquakeNPP: a benchmarking platform that curates and standardizes existing public resources: globally available earthquake catalogs, the ETAS model, and evaluation protocols from the seismology community. The datasets cover a range of small to large target regions within California, dating from 1971 to 2021, and include different methodologies for dataset generation. Benchmarking experiments, using both log-likelihood and generative evaluation metrics widely recognised in seismology, show that none of the five NPPs tested outperform ETAS. These findings suggest that current NPP implementations are not yet suitable for practical earthquake forecasting. Nonetheless, EarthquakeNPP provides a platform to foster future collaboration between the seismology and machine learning communities.
@@ 1 Introduction
Operational earthquake forecasting by global governmental organisations such as the US Geological Survey (USGS) necessitates the development of models which can forecast the times and locations of damaging earthquakes. While model development is ongoing in the seismology community, recent progress has relied upon refinement of a spatio-temporal point process model known as the Epidemic-Type Aftershock Sequence (ETAS) model (Ogata, 1988; 1998). This continued reliance on a low-dimensional parametric framework stands in contrast to the substantial growth in available earthquake data (Takanami et al., 2003; Shelly, 2017; Ross et al., 2019; White et al., 2019; Mousavi et al., 2020; Tan et al., 2021; Mousavi and Beroza, 2023).
In contrast, the machine learning community has offered promising advancements over classical point process models like ETAS with Neural Point Process (NPP) models, showcasing greater flexibility (Du et al., 2016; Omi et al., 2019a; Shchur et al., 2019; Jia and Benson, 2019; Chen et al., 2021; Zhou et al., 2022; Zhou and Yu, 2024). While some initial benchmarking of these models has been conducted on an earthquake dataset in Japan, these experiments lack relevance for stakeholders in the seismology community. The benchmark omits the largest earthquake sequence from the region, introduces data leakage with non-sequential train-test splits, and does not compare against state-of-the-art models like ETAS.
@@
Although initial benchmarking shows that none of the five tested NPP implementations outperform ETAS, EarthquakeNPP is designed to support ongoing model development and evaluation. In addition to the standard log-likelihood metric common in the NPP literature, the platform incorporates the generative evaluation procedures used in seismology for more rigorous benchmarking. This ensures that future NPPs (and other models such as time series approaches (Wang et al., 2017) and Bayesian point processes (Serafini et al., 2023)) can have direct relevance to seismological stakeholders. All datasets, experiments, and documentation are available at https://github.com/ss15859/EarthquakeNPP.
@@ 1.1.1 Benchmarking by the NPP Community
Chen et al. (2021) introduced an earthquake dataset for benchmarking the Neural Spatio-temporal Point Process (NSTPP) model using a global dataset from the U.S. Geological Survey, focusing on Japan from 1990 to 2020. They considered earthquakes with magnitudes above 2.5, splitting the data into month-long segments with a 7-day offset. They exclude earthquakes from November 2010 to December 2011, deeming these sequences "too long" and "outliers". However, this period includes the 2011 Tohoku earthquake (Mori et al., 2011), the largest earthquake recorded in Japan and the fourth largest in the world, at magnitude 9.0. This exclusion renders the benchmarking experiment irrelevant for seismologists, as it is precisely these large earthquakes and their aftershocks that are crucial to forecast due to their damaging impact.
@@
The dataset is partitioned into training, testing, and validation segments. Rather than following a chronological split that would reflect operational forecasting, the segments are assigned in an alternating pattern. This design introduces data leakage, as it misrepresents a realistic forecasting setup and artificially inflates performance measures due to the nature of earthquake triggering (Freed, 2005). Specifically, because the model is evaluated on windows that immediately precede its training windows, it can exploit backward-in-time causal dependencies. Section B.2 quantifies the resulting performance inflation, expressed in terms of information gain.
@@
Although earthquakes with magnitudes above 2.5 are considered by Chen et al. (2021), following a change in USGS policy on global data collection, from 2009 onwards, only events above magnitude 4.0 are recorded in the dataset. For earthquake forecasting in Japan, seismologists use datasets from Japanese data centers since they are more comprehensive and complete than global datasets. Section A.2 describes the biases incurred from such data missingness.
@@
Chen et al. (2021) benchmark their model against another spatio-temporal model, Neural Jump SDEs (Jia and Benson, 2019), and a temporal-only Hawkes process, even though a spatio-temporal Hawkes process would provide a more rigorous benchmark. Subsequent papers adopting this benchmark (Zhou et al., 2022; Yuan et al., 2023; Zhou and Yu, 2024) similarly lack comparisons to a spatio-temporal Hawkes process, benchmarking instead against temporal-only or spatial-only baselines or other spatio-temporal NPPs.
1.1.2 Benchmarking by the Seismology Community.
@@
Model comparison has been crucial in the development of earthquake forecasting models since their inception (Kagan and Knopoff, 1987; Ogata, 1988). The Collaboratory for the Study of Earthquake Predictability (CSEP) (Michael and Werner, 2018; Schorlemmer et al., 2018; Savran et al., 2022; Iturrieta et al., 2024) (https://cseptesting.org/ ) aims to unify the framework for earthquake model testing and evaluation, hosting retrospective and fully prospective forecasting experiments globally. CSEP benchmarks short-term models using performance metrics that require forecasts to be generated by simulating many repeat sequences over a specified time horizon (typically one day). These simulated forecasts are compared by discretizing time and space intervals, with test statistics calculated for event counts, magnitudes, locations, and times. The simulation-based approach allows the inclusion of generative models that do not output explicit earthquake probabilities (i.e., a likelihood), and enables evaluation of the full distribution of entire sampled sequences.
@@
Two existing works benchmark NPPs for earthquake forecasting within the seismology community. The first by Dascher-Cousineau et al. (2023) extends a temporal-only NPP from Shchur et al. (2019) to include earthquake magnitudes. The second by Stockman et al. (2023) extends another temporal-only model by Omi et al. (2019a) to target larger magnitude events. Both models are benchmarked against a temporal ETAS model, showing moderate improvements over the baseline. Extending these models to include spatial data is necessary for further testing and potential operational use in the seismological community.
@@ 2.1 Spatio-Temporal Point Processes
A spatio-temporal point process is a continuous-time stochastic process that models the random number of events $N(S\times(t_{a},t_{b}])$ which occur in a space-time interval $\mathcal{S}\times(t_{a},t_{b}],\ \mathcal{S}\subseteq\mathbb{R}^{2},\ (t_{a},t_{b}]\subset\mathbb{R}^{+}$. This process is typically defined by a non-negative conditional intensity function
@@
Most models specify the conditional intensity function, though some (e.g. Shchur et al., 2019; Chen et al., 2021; Yuan et al., 2023) directly model this pdf. Model parameters are typically estimated by maximizing the log-likelihood of observed events within a training time interval $T_{0},T_{1}]$ and spatial region $\mathcal{S}$,
@@
For model evaluation and comparison, the log-likelihood of observing events in the test set can be used as a performance metric. This is consistent with a wealth of literature in the seismology community (see Zechar et al., 2010, and references therein) as well as the wider general point process literature (Daley and Vere-Jones, 2004), which now includes neural point processes (Shchur et al., 2021). The metric evaluates models that output probability distributions over their predictions and consequently penalises models that are overconfident. Although evaluating on events in the test set, the test log-likelihood, $\log p\left((t_{i},\mathbf{x}_{i})|t_{i}\in[T_{2},T_{3}],\mathcal{H}_{T_{2}}\right)$, may still contain dependence upon events prior to the test window $T_{2},T_{3}]$, typically contained in the history $\mathcal{H}_{T_{2}}$ of the intensity function. Comparing the difference in mean log-likelihood per event provides the information gain from one model to another (Daley and Vere-Jones, 2004).
Point processes are the dominant modeling approach in the seismology community, used extensively in both real-time operational earthquake forecasting (Mizrahi et al., 2024a) and established benchmarking experiments (CSEP) (Taroni et al., 2018; Rhoades et al., 2018). The point process representation of earthquake data aligns naturally with their occurrence as discrete events in time (Kagan, 1994). Furthermore, this modeling approach is favored over discretized forecasting models (e.g., time series) because it eliminates the need for optimizing binning strategies and allows for immediate updates, rather than waiting until the end of a time bin - a delay that could miss critical, potentially damaging events.
@@ 2.2 ETAS
The Epidemic Type Aftershock Sequence (ETAS) model (Ogata, 1998) is a spatio-temporal Hawkes process Hawkes (1971); Siviero et al. (2024); Bernabeu and Mateu (2025) which models how earthquakes cluster in time and space. It has been adopted for operational earthquake forecasting by government agencies in California (Milner et al., 2020), New-Zealand (Christophersen et al., 2017), Italy (Spassiani et al., 2023), Japan (Omi et al., 2019b) and Switzerland (Mizrahi et al., 2024b), and performs consistently well in CSEP’s retrospective and fully prospective forecasting experiments (e.g. Woessner et al., 2011; Rhoades et al., 2018; Taroni et al., 2018; Cattania et al., 2018; Mancini et al., 2019; 2020; 2022). The general formulation of the model is
@@
where $\mu$ is a constant background rate of events, $g(\cdot,\cdot,\cdot)$ is a non-negative excitation kernel which describes how past events contribute to the likelihood of future events and $m_{i}$ are the associated magnitudes of each event. The equivalent formulation as a Hawkes branching process accompanies a causal branching structure $\mathbf{B}$. This concept broadly aligns with the understanding of the physics of earthquake triggering and interaction, e.g. via dynamic wave triggering (Brodsky and van der Elst, 2014) and static stress triggering (Gomberg, 2018; Mancini et al., 2020).
Although ETAS can be fit by maximizing the log-likelihood function directly, parameter estimation is typically performed by simultaneously estimating the branching structure $\mathbf{B}$. Veen and Schoenberg (2008) developed an Expectation Maximisation (EM) procedure, which maximises the marginal likelihood over the unobserved branching structure, $\log\int p(\mathcal{H}_{T_{1}}|\mathbf{B},\mathbf{\theta})p(\mathbf{B}|\mathbf{\theta})d\mathbf{B}$ through the iteration
@@
The formulation of the ETAS model we present in the EarthquakeNPP benchmark is implemented in the etas python package by Mizrahi et al. (2022). It defines the triggering kernel as
@@
where $r^{2}$ is the squared distance between events and $k,a,c,\omega,\tau,d,\gamma,\rho$ are the learnable parameters along with the constant background rate $\mu$. This triggering kernel is derived from statistical distributions found through decades of observational studies (Utsu and Seki, 1955; Utsu, 1970; Utsu et al., 1995) and several of the learnable parameters have been linked to physical properties of the earthquake rupture process (Utsu et al., 1995; Ide, 2013).
Despite its widespread use, it is commonly accepted that ETAS is a misspecified model of seismicity. By construction, ETAS describes only self-exciting triggering behaviour and therefore cannot capture inhibitory effects or stress relaxation processes, such as those represented by stress-release models (Zheng and Vere-Jones, 1991; Xiaogu and Vere-Jones, 1994; Bebbington and Harte, 2003), or by models based on elastostatic stress transfer and Coulomb rate-and-state friction (Dieterich, 1994). In addition, foreshock activity has been observed to deviate from ETAS assumptions, both spatially and temporally (McGuire et al., 2005; Brodsky, 2011; Lippiello et al., 2012; Ogata and Katsura, 2014). Finally, to simplify inference, ETAS typically assumes isotropic spatial triggering kernels, despite observational evidence for anisotropic and fault-aligned aftershock distributions (Page and van der Elst, 2022). Together, these limitations motivate the exploration of more flexible modelling frameworks capable of capturing richer spatio-temporal structure in earthquake sequences.
2.3 Neural Point Processes
Neural point processes (NPPs) have emerged in recent years within the machine learning literature as flexible alternatives to classical parametric point process models. Their central motivation is to replace restrictive, hand-crafted parametric forms with neural network based components that can learn complex, non-linear dependencies directly from data. This makes them particularly appealing for earthquake forecasting to overcome the known limitations of the ETAS model discussed in Section 2.2.
Early developments focused on temporal point processes (Shchur et al., 2021). Du et al. (2016) introduced the use of recurrent neural networks (RNNs) to encode the event history into a fixed-dimensional latent state, replacing explicit summation over past events with a learned representation of temporal dependence. Subsequent work explored alternative sequence encoders, including LSTMs (Mei and Eisner, 2017) and Transformers (Zuo et al., 2020), alongside a variety of decoding strategies for modelling the conditional intensity or inter-event time distribution (Du et al., 2016; Omi et al., 2019a; Shchur et al., 2019). In most cases, model parameters are learned by maximising the log-likelihood of observed event sequences, although alternative training objectives have also been proposed (Xiao et al., 2017; Li et al., 2018).
These temporal formulations were later extended to spatio-temporal settings (Mukherjee et al., 2025) by incorporating event locations into the history encoder and introducing flexible decoders for the conditional spatial distribution of future events. Existing spatio-temporal NPPs can be broadly grouped into three modelling classes. The first class extends Hawkes-type formulations by replacing parametric triggering kernels with neural network based influence functions, allowing non-stationary and history-dependent excitation (Zhou et al., 2022; Dong et al., 2022; Zhou and Yu, 2024). A second class models event dynamics in continuous time using neural ordinary differential equations, jointly evolving latent temporal states and spatial distributions (Jia and Benson, 2019; Chen et al., 2020). A third, more recent class adopts fully generative approaches based on diffusion or score matching, learning the joint spatio-temporal distribution of events without explicitly parameterising an intensity function (Yuan et al., 2023; Li et al., 2023; Lüdke et al., 2024). These approaches differ substantially in their computational cost, interpretability, and suitability for simulation and likelihood-based evaluation; see Mukherjee et al. (2025) for a detailed discussion.
Figure 1: Earthquakes contained in the observational datasets found in EarthquakeNPP. Colours indicate the respective datasets, including the target region, magnitude of completeness $M_{c}$, number of events and the time period that the dataset spans. In red is a fault map from the GEM Global Active Faults Database (Styron and Pagani, 2020).
@@ 3 EarthquakeNPP Datasets
The EarthquakeNPP datasets (Figure 1) encompass earthquake records, including timestamps, geographical coordinates, and magnitudes, documented within California from 1971 to 2021. California, with its dense network and high seismic hazard, has been extensively studied, demonstrating the utility of forecasting algorithms (Gerstenberger et al., 2004; Field, 2007; Field et al., 2021). It encompasses the San Andreas fault plate boundary system (Zoback et al., 1987) and includes modern high-resolution catalogs with numerous small magnitude earthquakes, offering potential for new, more expressive models.
@@
Development The U.S. Geological Survey (USGS) National Earthquake Information Center (NEIC) monitors global earthquakes (Mw 4.5 or larger) and provides complete seismic monitoring of the United States for all significant earthquakes (> Mw 3.0 or felt). Its contributing seismic networks have produced the Advanced National Seismic System (ANSS) Comprehensive Catalog of Earthquake Events and Products. The Southern California Seismic Network (SCSN) has developed and maintained the standard earthquake catalog for Southern California (Hutton et al., 2010) since the Caltech Seismological Laboratory began routine operations in 1932. Significant network improvements since the 1970s and 1980s reduced the catalog completeness from Mw 3.25 to Mw 1.8. White et al. (2019) created an enhanced catalog focusing on the San Jacinto fault region, using a dense seismic network in Southern California. This denser network, combined with automated phase picking (STA/LTA), ensures a 99% detection rate for earthquakes greater than Mw 0.6 in a specific subregion (White et al., 2019). Using data collected by the SCSN, Ross et al. (2019) generated a denser catalog by reanalyzing the same waveform data with a template matching procedure that looks for cross-correlations with the wavetrains of previously detected events.
@@
forecasting (USGS utilizes ComCat in aftershock forecasts they release to the public).
Three magnitude thresholds (Mw 2.0, 2.5, 3.0 for SCEDC_20, SCEDC_25, SCEDC_30) explore the effect of truncation on forecasting model performance. To explore if newly detected low magnitude earthquakes contain additional predictive information. To explore if newly detected low magnitude earthquakes contain additional predictive information (with different detection methodology).
A central challenge when working with earthquake catalogs is data missingness, referred to in seismology as catalog incompleteness (Mignan and Woessner, 2012). Earthquakes are assumed to be only fully detected above a time- and region-dependent completeness magnitude $M_{c}$, which reflects limitations of the seismic network and changes in detection capability over time. Ignoring this incompleteness can introduce substantial bias into both model fitting and evaluation (Sornette and Werner, 2005), particularly for methods that rely on smaller magnitude events.
All EarthquakeNPP datasets are constructed from publicly available raw catalogs provided by their respective data centres. To enable a consistent and realistic retrospective forecasting experiment, the raw data is preprocessed by restricting each dataset to a target spatial region, truncating events above a dataset-specific magnitude threshold $M_{\text{cut}}\geq M_{c}$ (e.g., Mignan et al., 2011; Mignan and Woessner, 2012) and removing duplicate locations.
Notebooks to access and preprocess the datasets, along with the associated benchmarking experiments, are publicly available at https://github.com/ss15859/EarthquakeNPP, accompanied by detailed dataset documentation. A more in-depth discussion of earthquake catalog generation, completeness, and preprocessing choices is provided in Appendix A. Table 1 summarises the key characteristics of each EarthquakeNPP dataset.